OpenStudy (amber15):

please Help !! The intensity I of light bulb measured in watts per square meter (w/m^2), varies inversely as the square of the distance d from the light bulb . . suppose I is 55 w/m^2 when the distance is 8 m . find the intensity when the distance from the light bulb is 5 m away

OpenStudy (anonymous):

The first sentence tells you \[\text{intensity}=\frac{\text{power}}{(\text{distance})^2}~~\iff~~I=\frac{P}{d^2}\] Given that \(I=55\) when \(d=8\), you can find the corresponding value of \(P\): \[55=\frac{P}{8^2}~~\implies~~P=\cdots\] If the power \(P\) is kept constant, then the intensity \(I\) will still change as a function of the distance \(d\). Using this \(P\), you would then find \(I\) when \(d=5\) by solving \[I=\frac{P}{5^2}~~\implies~~I=\cdots\]

OpenStudy (amber15):

oh ok